OpenStudy (anonymous):

The size of a bicycle is determined by the diameter of the wheel. You want a 20-in bicycle for your birthday. What is the area of one of the wheels? Answer 385.47 in2 314.16 in2 484.37 in2 345.54 in2

OpenStudy (anonymous):

@cutepochacco

OpenStudy (emmigrace222):

let me do this :)

OpenStudy (emmigrace222):

ok so do you know the steps you need to do this?

OpenStudy (anonymous):

diameter=12 inches , so the radius will be 6 inches Area for a circle is \(A=\pi r^2\), substitute the value of your radius in the formula.

OpenStudy (anonymous):

what

OpenStudy (emmigrace222):

ok wolf you plug in 6 witch is the radius and 12 for pie I believe...

OpenStudy (emmigrace222):

\[6\times6=36\]

OpenStudy (emmigrace222):

then take 12 and times it right @Data_LG2 ?

OpenStudy (anonymous):

12?????! o.0

OpenStudy (emmigrace222):

its been a while sense i did this kind or problem i forgot what to do with the diameter! lol

OpenStudy (emmigrace222):

im trying my best though

OpenStudy (anonymous):

 \(\LARGE DIAMETER=\frac{1}{2}RADIUS\) \(\LARGE AREA\ OF\ CIRCLE= \pi r^2\), where r=radius why I am using area of the circle to calculate the area of a wheel? because I believe that a wheel is a circle, don't you think?

OpenStudy (anonymous):

You are given the value of the \(\Large DIAMETER\), which is \(\Large 12 inches\). To get the radius, divide the diameter by 2. THEN plug the value of the radius into the formula.

OpenStudy (emmigrace222):

well how do you get pie though

OpenStudy (anonymous):

\(\LARGE \pi=3.145\) , this is a FAMOUS constant!

OpenStudy (emmigrace222):

ohhhhhhh!!

OpenStudy (anonymous):

understand?

OpenStudy (emmigrace222):

i do now i remember 100% now

OpenStudy (anonymous):

glad to know :)

OpenStudy (anonymous):

you must remember all kind of things because I can assure you that it will really affect you and may help you someday . ^_^